OpenStudy (anonymous):

? Which of the following is the simplified form of 4 times x cubed over 9 times y to the fourth power divided by 8 times x to the fifth power over 3 times y squared ? 6x2y2 6 over x squared times y squared 1 over 6 times x squared times y squared x squared times y squared over 6

OpenStudy (campbell_st):

so you are looking at \[\frac{\frac{4x^3}{9y^4}}{\frac{8x^5}{3y^2}}\] the rule is flip the denominator and multiply so you are looking at \[\frac{4x^3}{9y^4} \times \frac{3y^2}{8x^5}\] cancel the common factors, whats left is the answer.

OpenStudy (anonymous):

None of them are common.

OpenStudy (campbell_st):

really well lets look at it bits \[\frac{12x^3y^2}{72x^5y^4} = \frac{12}{72} \times \frac{x^3}{x^5} \times \frac{y^2}{y^4}\] can you simplify anything the the expanded form above.

OpenStudy (anonymous):

>.<

OpenStudy (campbell_st):

what does 12/72 simplify to...?

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

or its .16

OpenStudy (campbell_st):

not quite.... 12/72 = 1/6 divide the numerator and denominator by 12 and you need a fraction answer... next... \[\frac{x^3}{x^5} = \frac{x \times x \times x}{x \times x \times x \times x \times x}\] what do you think this will simplify to...?